A secretary typed 59 words per minute for 59 minutes, 42 words per minutes for 42 minutes and 73 words per minute for a 100 minutes. What was the average rate of her typing speed in words per minute for this time?
Answer by Theo(13342)

The total minutes she typed is 59 + 42 + 100 = 201 minutes

The total number of words she typed is 59*59 + 42*42 + 73*100 = 12545 words

The average overall speed is 12545 / 201 = 62.41293532 words per minute.